Let X be normally distributed with mean μ. 126 and standard deviation σ-22. [You may find it useful to reference the ztable.] a. Find AX s 100). (Round value to 2 decimal places and final answer to 4 decimal places.) POX S 100) b. Find P95 sX s110) (Roundr value to 2 decimal places and final answer to 4 decimal places.) P(95 SXS 110) c. Find x such that PXsX)-0.410. (Round value and final answer to 3 decimal places.) d. Find x such that RX >X) = 0.820. (Round z value and final answer to 3 decimal places.)
Answer #1

Solution:

Given μ=126 and σ=22

a)

we have P( x <= 100) = P( x < = x - μ /  μ)= ( 99.5 -126 /22 ) = -1.204

P ( X <= -1.204)

= P( Z < -1.204)

=1−P ( Z<1.204 )

=1−0.8849

=0.1151

b)  Given that P( 95 ≤ X  ≤ 100)

P( 95 ≤ X  ≤ 100) = P( 94.5 - 126 /22  ≤ X  ≤ 100 -126/ 22)

= ( -1.431  ≤ X  ≤ -1.181 )

P( -1.431  ≤ X  ≤ -1.181 ) = P( -1.431  ≤ Z  ≤ -1.181 )

= 0.0426 ( by standard normal table)

c) given that P(X ≤ x) =0.410

x =μ - Z* σ = 126 - ( 0.2275 *22) = 120.995

X = 120.995 -0.5 = 121.495

d) given that P(X ≤ x) =0.820

x =μ - Z* σ = 126 + ( .9513*22) = 146.137